# Mastering Text-to-Image Diffusion: Recaptioning, Planning, and Generating with Multimodal LLMs

### v2

@@ 1 Introduction
Recent advancements in diffusion models (Sohl-Dickstein et al., 2015; Dhariwal & Nichol, 2021; Song et al., 2020; Yang et al., 2023c) have significantly improve the synthesis results of text-to-image models, such as Imagen (Saharia et al., 2022), DALL-E 2/3 (Ramesh et al., 2022; Betker et al., 2023) and SDXL (Podell et al., 2023). However, despite their remarkable capabilities in synthesizing realistic images consistent with text prompts, most diffusion models usually struggle to accurately follow some complex prompts (Feng et al., 2022; Lian et al., 2023; Liu et al., 2022; Bar-Tal et al., 2023), which require the model to compose objects with different attributes and relationships into a single image (Huang et al., 2023a).
@@
Some works begin to solve this problem by introducing additional layouts/boxes (Li et al., 2023b; Xie et al., 2023; Yang et al., 2023e; Qu et al., 2023; Chen et al., 2024; Wu et al., 2023b; Lian et al., 2023) as conditions or leveraging prompt-aware attention guidance (Feng et al., 2022; Chefer et al., 2023; Wang et al., 2023) to improve compositional text-to-image synthesis. For example, StructureDiffusion (Feng et al., 2022) incorporates linguistic structures into the guided generation process by manipulating cross-attention maps in diffusion models. GLIGEN (Li et al., 2023b) designs trainable gated self-attention layers to incorporate spatial inputs, such as bounding boxes, while freezing the weights of original diffusion model.
@@ 2.2 Text-to-image Generation
1.Task instruction You are an smart region planner for image. You should use split ratio to specify the split method of the image, and then recaption each subregion prompts with more descriptive prompts while maintaining the original meaning. 2.Multi-modal split tutorial …… 3. In-context examples User Prompt: A girl with white ponytail and black dress are chatting with a blonde curly hair girl in a white dress in a cafe. # Key pharses extraction and Recaption # Split ratio Planning # Composition Logic # Aesthetic Considerations: # Final output 4.Trigger CoT reasoning ability of MLLMs User Prompt: An old man with his dog is looking at a parrot on the tree. Reasoning: Let’s think step by step……
To facilitating inferring the region for each subprompt, we adhere to three key principles in designing in-context example and generating informative rationales: (i) the objects with same class name (e.g., five apples) will be separately assign to different regions to ensure the numeric accuracy; (ii) for complex interactions between two entities, we take these two entities as a whole to avoid contradictory overlapped generation results mentioned in (Lian et al., 2023); (iii) If the prompt focuses more on the appearance of a specific entity, we treat the different parts of this entity as different entities (e.g., A green hair twintail in red blouse , wearing blue skirt. $\Longrightarrow$ green hair twintail, red blouse, blue skirt).
@@ 2.3 Text-Guided Image Editing
Model Attribute Binding Object Relationship Complexnormal-$\uparrow$
Color normal-$\uparrow$ Shapenormal-$\uparrow$ Texturenormal-$\uparrow$ Spatialnormal-$\uparrow$ Non-Spatialnormal-$\uparrow$
@@ 3.1 Text-to-Image Generation
Our RPG is general and extensible, we can incorporate arbitrary MLLM architectures and diffusion backbones into the framework. In our experiment, we choose GPT-4 (OpenAI, 2023) as the recaptioner and CoT planner, and use SDXL (Podell et al., 2023) as the base diffusion backbone to build our RPG framework. Concretely, in order to trigger the CoT planning ability of MLLMs, we carefully design task-aware template and high-quality in-context examples to conduct few-shot prompting. Base prompt and its weighted hyperparameter base ratio are critical in our regional diffusion, we have provide further analysis in Figure 16. When the user prompt includes the entities with same class (e.g., two women, four boys), we need to set higher base ratio to highlight these distinct identities. On the contrary, when user prompt includes the the entities with different class name (e.g., ceramic vase and glass vase), we need lower base ratio to avoid the confusion between the base prompt and subprompts.
@@ 5 Related Work
Diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Song & Ermon, 2020; Song et al., 2020) are a promising class of generative models, and Dhariwal & Nichol (2021) have demonstrated the superior image synthesis quality of diffusion model over generative adversarial networks (GANs) (Reed et al., 2016; Creswell et al., 2018). GLIDE (Nichol et al., 2021) and Imagen (Saharia et al., 2022) focus on the text-guided image synthesis, leveraging pre-trained CLIP model (Radford et al., 2021; Raffel et al., 2020) in the image sampling process to improve the semantic alignment between text prompt and generated image. Latent Diffusion Models (LDMs) (Rombach et al., 2022) move the diffusion process from pixel space to latent space for balancing algorithm efficiency and image quality. Recent advancements in text-to-image diffusion models , such as SDXL (Podell et al., 2023) Dreambooth (Ruiz et al., 2023) and DALL-E 3 (Betker et al., 2023), further improve both quality and alignment from different perspectives. Despite their tremendous success, generating high-fidelity images with complex prompt is still challenging (Ramesh et al., 2022; Betker et al., 2023; Huang et al., 2023a). This problem is exacerbated when dealing with compositional descriptions involving spatial relationships, attribute binding and numeric awareness. In this paper, we aim to address this issue by incorporating the powerful CoT reasoning ability of MLLMs into text-to-image diffusion models.

### v3

@@ 1 Introduction
Recent advancements in diffusion models (Sohl-Dickstein et al., 2015; Dhariwal & Nichol, 2021; Song et al., 2020; Yang et al., 2023c) have significantly improve the synthesis results of text-to-image models, such as Imagen (Saharia et al., 2022), DALL-E 2/3 (Ramesh et al., 2022; Betker et al., 2023) and SDXL (Podell et al., 2023). However, despite their remarkable capabilities in synthesizing realistic images consistent with text prompts, most diffusion models usually struggle to accurately follow some complex prompts (Feng et al., 2022; Lian et al., 2023; Liu et al., 2022; Bar-Tal et al., 2023), which require the model to compose objects with different attributes and relationships into a single image.
@@
Some works introduce additional layouts/boxes (Li et al., 2023b; Xie et al., 2023; Yang et al., 2023e; Qu et al., 2023; Chen et al., 2024; Wu et al., 2023b; Lian et al., 2023) as conditions or leveraging prompt-aware attention guidance (Feng et al., 2022; Chefer et al., 2023; Wang et al., 2023) to improve compositional text-to-image synthesis. For example, StructureDiffusion (Feng et al., 2022) incorporates linguistic structures into the guided generation process by manipulating cross-attention maps in diffusion models. GLIGEN (Li et al., 2023b) designs trainable gated self-attention layers to incorporate spatial inputs, such as bounding boxes, while freezing the weights of original diffusion model.
@@ 2.2 Text-to-image Generation
1.Task instruction
You are an smart region planner for image. You should use split ratio to specify the split method of the image, and then recaption each subregion prompts with more descriptive prompts while maintaining the original meaning.
2.Multi-modal split tutorial
…… 3. In-context examples
User Prompt: A girl with white ponytail and black dress are chatting with a blonde curly hair girl in a white dress in a cafe.
# Key pharses extraction and Recaption
# Split ratio Planning
# Composition Logic
# Aesthetic Considerations:
# Final output
4.Trigger CoT reasoning ability of MLLMs
User Prompt: An old man with his dog is looking at a parrot on the tree.
Reasoning: Let’s think step by step……
To facilitating inferring the region for each subprompt, we adhere to two key principles in designing in-context example and generating informative rationales: (i) the objects with same class name (e.g., five apples) will be separately assign to different regions to ensure the numeric accuracy; (ii) If the prompt focuses more on the appearance of a specific entity, we treat the different parts of this entity as different entities (e.g., A green hair twintail in red blouse , wearing blue skirt. $\Longrightarrow$ green hair twintail, red blouse, blue skirt).
@@ 2.3 Text-Guided Image Editing
[…]
@@ 3.1 Text-to-Image Generation
Our RPG is general and extensible, we can incorporate arbitrary MLLM architectures and diffusion backbones  [^1] https://github.com/CompVis/stable-diffusion [^2] https://github.com/huggingface/diffusers [^3] https://github.com/hako-mikan/sd-webui-regional-prompter into the framework. In our experiment, we choose GPT-4 (OpenAI, 2023) as the recaptioner and CoT planner, and use SDXL (Podell et al., 2023) as the base diffusion backbone to build our RPG framework. Concretely, in order to trigger the CoT planning ability of MLLMs, we carefully design task-aware template and high-quality in-context examples to conduct few-shot prompting. Base prompt and its weighted hyperparameter base ratio are critical in our regional diffusion, we have provide further analysis in Figure 16. When the user prompt includes the entities with same class (e.g., two women, four boys), we need to set higher base ratio to highlight these distinct identities. On the contrary, when user prompt includes the the entities with different class name (e.g., ceramic vase and glass vase), we need lower base ratio to avoid the confusion between the base prompt and subprompts.
@@ 5 Related Work
Diffusion models (Sohl-Dickstein et al., 2015; Song & Ermon, 2019; Ho et al., 2020; Song & Ermon, 2020; Song et al., 2020; Yang et al., 2024a) are a promising class of generative models, and Dhariwal & Nichol (2021) have demonstrated the superior image synthesis quality of diffusion model over generative adversarial networks (GANs) (Reed et al., 2016; Creswell et al., 2018). GLIDE (Nichol et al., 2021) and Imagen (Saharia et al., 2022) focus on the text-guided image synthesis, leveraging pre-trained CLIP model (Radford et al., 2021; Raffel et al., 2020) in the image sampling process to improve the semantic alignment between text prompt and generated image. Latent Diffusion Models (LDMs) (Rombach et al., 2022) move the diffusion process from pixel space to latent space for balancing algorithm efficiency and image quality. Recent advancements in text-to-image diffusion models , such as SDXL (Podell et al., 2023), ContextDiff (Yang et al., 2024b) and DALL-E 3 (Betker et al., 2023), further improve both quality and alignment from different perspectives. Despite their tremendous success, generating high-fidelity images with complex prompt is still challenging (Betker et al., 2023; Huang et al., 2023a). This problem is exacerbated when dealing with compositional descriptions involving spatial relationships, attribute binding and numeric awareness. In this paper, we aim to address this issue by incorporating the powerful CoT reasoning ability of MLLMs into text-to-image diffusion models.